### v3

# “Cold, Calculated, and Condescending”: How AI Identifies and Explains Ableism Compared to Disabled People

Abstract.
People with disabilities (PwD) regularly encounter ableist hate and microaggressions online. These spaces are generally moderated by machine learning models, but little is known about how effectively AI models identify ableist speech and how well their judgments align with PwD. To investigate this, we curated a first-of-its-kind dataset of 200 social media comments targeted towards PwD, and prompted state-of-the art AI models (i.e., Toxicity Classifiers, LLMs) to score toxicity and ableism for each comment, and explain their reasoning. Then, we recruited 190 participants to similarly rate and explain the harm, and evaluate LLM explanations. Our mixed-methods analysis highlighted a major disconnect: AI underestimated toxicity compared to PwD ratings, while its ableism assessments were sporadic and varied. Although LLMs identified some biases, its explanations were flawed—they lacked nuance, made incorrect assumptions, and appeared judgmental instead of educational. Going forward, we discuss challenges and opportunities in designing moderation systems for ableism, and advocate for the involvement of intersectional disabled perspectives in AI.
large language models, disability, ableism, natural language, toxicity classifiers, hate speech
journalyear: 2025copyright: acmlicensedconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732128isbn: 979-8-4007-1482-5/2025/06ccs: Social and professional topics People with disabilitiesccs: Computing methodologies Artificial intelligence
Figure 1. A representation of AI models being unable to identify and explain ableism, as it relates to lived disability experiences.
[…]
A diagram comparing how different evaluators: Large Language Models (LLMs), Toxicity Classifiers (TCs), People with disabilities (PwD), and People without disabilities—rate and explain harm from a dataset comment: ’It’s survival of the fittest.’ The flow begins with ’Dataset Comments’ on the left, showing colored dots representing various comments, leading to four evaluator categories. On the right, there are different harm measurements. PwD assess the comment as ’Not ableist, but toxic,’ with a 20% harm rating, explaining that ’Fetishizing people is always wrong.’ TCs provide a 38% toxicity score. LLMs rate the comment as 60% harmful, stating ’This is significantly ableist.’
Content Warning: This paper contains graphic examples of explicit, offensive, and ableist language.
1. Introduction
An estimated 1.3 billion people worldwide experience a significant disability (Organization, 2022), a number that only partially represents all people that experience disability from aging, trauma, injury, or other impediments. People with disabilities (PwD) routinely experience violence, victimization, and discrimination at nearly four times the rate than people without disabilities (Non-PwD) (Harrell, 2021). These experiences inevitably bleed into online spaces through hostile language, such as microaggressive comments (Xiong, 2016; Heung et al., 2022; Keller and Galgay, 2010; Sannon et al., 2023), or explicit mentions of violence and aggression (Lyu et al., 2024; Heung et al., 2024; Alhaboby et al., 2016; Burch, 2018; Ringland, 2019).
Online platforms rely on AI models like toxicity classifiers (TCs) and large language models (LLMs) to combat online hate. These models demonstrated success in moderating vast online spaces (Jain et al., 2018; Muralikumar et al., 2023; Friedl, 2023), especially more recent LLMs that show promise in moderating a range of harmful speech (Kumar et al., 2024; Franco et al., 2023). However, there is a growing concern that these models replicate existing societal biases and mirror harmful behavior against historically marginalized groups (Blodgett and O’Connor, 2017; Franco et al., 2023; Hartvigsen et al., 2022; Röttger et al., 2021; Halevy et al., 2021; Kumar et al., 2021; Sun et al., 2019). We see this harm surface in false moderation of speakers with specific dialects (Rios, 2020; Halevy et al., 2021; Elnagar et al., 2021; Sap et al., 2019), or in misclassification of queer vocabulary (Dias Oliva et al., 2021; Dorn et al., 2024). In these hostile environments, it is unclear how these models serve and represent the needs of disabled people. As it stands, the marginalization of PwD permeates into AI models and no study has examined how these models identify toxic and ableist speech targeted towards PwD.
Ableist speech is contextual and nuanced, often rooted in everyday language (e.g, “blind review process”) (Nikki, 2022), do AI models classify these phrases as ableist? Are they able to explain their predictions? Classification and reasoning of harm is not only increasingly important in emerging AI-assisted moderation practices (Mathew et al., 2022; Liu et al., 2019; Gillespie, 2019), but laws such as General Data Protection Regulation (GDPR) (Parliament and of the Council, 2016) in Europe have recently established a “right to explanation,” for decisions made based on AI applications. To establish safe and inclusive online spaces for disabled people, there is an urgent need to assess alignment between PwD and machine learning moderation systems. We pose the following research questions:
  1. RQ1:
    How well do toxicity classifiers and large language models identify ableist comments compared to PwD?
  2. RQ2:
    How well do large language models explain their ableist speech classification compared to PwD?
To address these questions, we first curated a labeled dataset of 200 social media comments, comprising of 175 ableist and 25 non-ableist comments. We prompted state-of-the-art TCs (PerspectiveAPI, AzureAI Content Safety API, OpenAI Content Moderation API) and LLMs (GPT-4, Gemini, Claude, Llama) to rate toxicity and ableism levels for each comment, asking LLMs to provide justifications for their ratings. Then, we recruited 190 participants to perform the same task—rate and explain ableist harm—and to evaluate a subset of explanations generated by the LLMs. Using a mixed-methods approach, we analyzed how closely assessments by PwD aligned with AI models.
Our findings revealed a clear misalignment between AI models and PwD in rating toxicity and ableism. Toxicity classifiers and LLMs (Gemini, GPT-4, and Claude) rated toxicity significantly lower than PwD, whereas ableism ratings by LLMs were varied. Notably, GPT-4, Claude, and Llama overrated ableism compared to PwD, especially for non-ableist comments. These inconsistencies raise concerns about the reliability of AI models in identifying ableist content. Explanations by LLMs were also misaligned and fell short of expectations of PwD. While PwD appreciated that LLMs identified common ableist aggressions and explained ableism succinctly, they found LLM explanations to be generic, clinical, and judgmental. They disagreed with incorrect and presumptuous claims made by LLMs, emphasizing that ableism explanations should educate the reader and build empathy for lived experiences of PwD.
Our results highlight that SOTA models do not account for toxic speech targeted towards PwD, thereby leaving the internet unsafe for one-sixth of the global population. Our focus on ableism explanations emphasizes how biased AI assessments could misrepresent disability narratives, making tone, language, and verbiage critical in addressing disability bias. We contribute the first publicly available dataset of ableist speech (see link), and collaborate with PwD to assess how AI systems identify and explain disability bias, setting a pivotal foundation to bridging this gap. We discuss the challenges of benchmarking ableist speech, and provide concrete avenues for online ableism harm interventions. This effort calls for the active involvement of PwD in the AI development process, ensuring that disabled experiences are represented and prioritized.
2. Related Work
2.1. Identifying and Moderating Harmful Speech Online
Online platforms use moderation to remove and reduce harmful content, such as derogatory language (Bilewicz and Soral, 2020; Castaño-Pulgarín et al., 2021), misinformation (Fernandez and Alani, 2018; Swire-Thompson and Lazer, 2020; Del Vicario et al., 2016), or even threats of violence (Olteanu et al., 2018; Mathew et al., 2019). Although some moderation is done via community volunteers (Gillespie, 2019), the sheer volume of emerging content has driven platforms to adopt automated moderation approaches. These approaches often rely on language models such as toxicity classifiers (TCs) to detect harmful speech (API, 2024; Kumar et al., 2021; Fortuna et al., 2020). Among these models, Google Jigsaw’s Perspective has emerged as a state-of-the-art toxicity classifier, trained on over 63 million Wikipedia comments (Friedl, 2023), and handling over 500 million requests per day to moderate content across platforms like Reddit, New York Times, OpenWeb, and Disqus (API, 2024). Due to its widespread use, several researchers have extensively audited and evaluated its performance (Kumar et al., 2021; Jain et al., 2018; Muralikumar et al., 2023), such as work of Muralikumar et al. (2023) on assessing the alignment between human toxicity ratings and Perspective’s ratings. More recently, with the rise of large language models (LLMs), many researchers have evaluated how well these models identify toxic content online (Franco et al., 2023; Kumar et al., 2024). For example, Kumar et al. (2024) evaluated OpenAI’s GPT-4 model and found it to be equally proficient at toxicity detection, achieving a median precision of 83% on selected subreddits.
While most online platforms rely on AI advances to identify toxic content, an emerging body of research shows that TCs and language models perpetuate societal biases and led to further discrimination and marginalization of users with marginalized identities of gender, race, and other attributes (Sap et al., 2019, 2022; Dias Oliva et al., 2021; Lyu et al., 2024; Heung et al., 2024). For example, language models have been shown to misclassify regional dialects, such as African-American Vernacular English, as toxic (Sap et al., 2019; Halevy et al., 2021; Elnagar et al., 2021), or mislabel and suppress vocabulary popular in queer communities (Dias Oliva et al., 2021; Dorn et al., 2024). These harms stem in part from who annotates the training data and whose voices are represented in toxicity benchmarking datasets (Blodgett and O’Connor, 2017; Blodgett et al., 2020; Selvam et al., 2023; Rauh et al., 2022). Annotator identities play a significant role in shaping how models interpret toxicity (Goyal et al., 2022; Kumar et al., 2021; Garg et al., 2023; Cabitza et al., 2023), and performance has shown improvement when annotations from historically marginalized groups were incorporated. For example, Goyal et al. (Goyal et al., 2022) demonstrated that creating “specialized rater pools” of African American or Queer raters enhanced model outcomes.
While substantial progress has been made in identifying racial and gender biases in toxicity detection, little is known about how these models perform on toxic speech targeted towards individuals with disabilities–a population that is facing increasing amounts of hate and harassment (Ringland, 2019; Kaur et al., 2024; Sannon et al., 2019; Heung et al., 2024; Sannon et al., 2023). These systems not only promise sensitive content moderation, but plan to inform content policy guidelines (Weng et al., 2023), exacerbating the need to involve disabled perspectives in moderating harm. Our work addresses this critical gap by curating the first-ever dataset dedicated to ableist speech, and examining alignment between AI models and people with disabilities. By actively involving disabled users in evaluating AI performance, we take a pivotal step toward addressing online ableism and advocating for safer, more inclusive digital spaces.
2.2. AI and Ableism
People with disabilities (PwD) increasingly use online platforms to advocate for disability rights and challenge regressive ableist norms (Ferguson, 216; Mann, 2018; Auxier et al., 2019). However, this growing advocacy has led to increased incidents of hate and harassment towards them (Ringland, 2019; Kaur et al., 2024; Sannon et al., 2019; Heung et al., 2024; Sannon et al., 2023). Recent research has documented the discrimination and toxic speech that PwD encounter online. From facing overt forms of hate (e.g., slurs, threats) (Sannon et al., 2023; Heung et al., 2024) to disguised hurtful language (e.g., invasive questions, infantilizing comments, denial of identity) (Keller and Galgay, 2010; Heung et al., 2022; Olkin et al., 2019; Lyu et al., 2024), many PwD feel excluded, bullied, and abused online (Burch, 2018), and several have forgoed their online communities (Heung et al., 2024; Lyu et al., 2024). Not only do online platforms fail to prevent toxic speech, they often suppress advocacy content posted by PwD (Heung et al., 2024; Sannon et al., 2023). Furthermore, there is a growing concern that machine learning models reinforce and perpetuate biases against PwD (Sheng et al., 2021; Hutchinson et al., 2020; Hassan et al., 2021; Gadiraju et al., 2023; Nakamura, 2019). Research shows that language models reinforce ableist stereotypes in text completion, make assumptions about PwD wanting to be “fixed,” and shift sentiments from positive to negative when disability-related terms are introduced in text  (Hassan et al., 2021; Gadiraju et al., 2023; Venkit et al., 2022; Whittaker et al., 2019).
In efforts to mitigate these biases, scholars have raised important questions on disability fairness in emerging AI technologies. As  Whittaker et al. (2019) and Trewin (2018) observe, the fundamentally diverse and nuanced nature of disability often makes it an “outlier” in machine learning datasets, with such outliers frequently treated as noise and disregarded. Disability is often “implicitly understood to be undesirable,” and AI is positioned to “solve the ‘problem’ of disability,” for example, teaching children with autism to act more neurotypically (Spiel et al., 2019; Whittaker et al., 2019). Similarly, in online communities, PwD are pressured to conform to platform policies (e.g., for people with eating disorders (Feuston et al., 2020)) and toxicity classifiers are designed to adhere to these policies (Yan et al., 2019; Zhou et al., 2020).
While prior research has emphasized the importance of explaining outputs of language models designed to identify toxicity  (Gao et al., 2017; Yadav et al., 2024; Han and Tsvetkov, 2020), there is a scarcity of research on how well state-of-the-art language models can identify and explain ableism—a crucial step needed to make online platforms safe for PwD  (Heung et al., 2022; Venkit et al., 2022; Keller and Galgay, 2010).
3. Methodology
In line with recent calls for mixed-method research in AI fairness assessments (Van Berkel et al., 2023), we adopted quantitative and qualitative approaches to evaluate alignment between AI systems and PwD.
Figure 2. A high-level overview of the study methodology.
[…]
The diagram illustrates a process where ableist and non-ableist comments are evaluated for harm by different groups. On the left is a table with three sample comments, their associated disabilities, and the intention behind the comments: 1.’Autism is so over-diagnosed these days’ — Disability: Autism Spectrum Disorder, Intention: Implicit (labeled with a red dot indicating ableism). 2. ’Life is not worth living if you can’t walk’ — Disability: Orthopedic/Physical Disability, Intention: Explicit (labeled with a red dot indicating ableism). 3. ’Do you need closed captioning?’ — Disability: d/Deaf or Hard of Hearing, Intention: Non-Ableist (labeled with a green dot). On the right, these comments are assessed for harm by four groups: Large Language Models (LLMs) (GPT-4, Gemini, Claude, Llama) with outputs going into ’Assessments of Harm.’ Toxicity Classifiers (Perspective, AzureAI, OpenAI) with outputs going into ’Assessments of Harm.’ People with disabilities with outputs forming the ’Ground Truth.’ People without disabilities with outputs also going into ’Assessments of Harm.’ The diagram represents how comments are evaluated for harm by both AI models and human groups.
3.1. Ableist Speech Dataset Curation
Gathering Ableist and Non-Ableist Data. Given the absence of publicly available datasets for ableist speech, we assembled our own dataset of ableist and non-ableist sentences. In line with prior investigations of discriminatory experiences on social media (Brock, 2012; McGregor and Mourão, 2016; Farrell et al., 2019; Yang and Counts, 2018), we used Twitter and Reddit to search for explicit mentions or call-outs to ableist speech by people with disabilities. We used hashtags and keywords such as ”offensive,” ”ableist,” ”hate,” ”harassment,” ”problematic,” and ”discrimination,” and manually filtered for posts that targeted users with disabilities. To represent a diverse range of disabilities, as put forth by the Individuals with Disabilities Education Act (IDEA) (of Education, 2003), we also searched for these keywords in specific subreddit communities, such as r/autism, r/ADHD, r/blind, r/physicaldisabilities, and r/deaf, among others. We limited our date range to 2019-2024 to capture more recent instances of ableism. In addition to social media data, we also gathered primary accounts by PwD as reported in academic papers on ableist speech (Sannon et al., 2023; Heung et al., 2022, 2024; Lyu et al., 2024).
We also gathered some non-ableist comments to assess for false positives, in line with benchmarking practices that require representations from both binaries of an attribute to detect disparities (e.g., ableist and non-ableist), for example, #DisabledOnIndiaTwitter (Mondal et al., 2022). We selected social media comments that were either unrelated to disability, (i.e., ”Reddit is so toxic sometimes!”), or comments that shared personal disability-related experiences, (i.e., “i can’t wear masks due to a cardiac issue”). In total, we collected 25 non-ableist comments in addition to 175 ableist comments, creating a comprehensive dataset of 200 comments.
Labeling Data. To ensure broad coverage in our dataset, we labelled all comments with (1) the specific disability it targeted (e.g., Blind or Low Vision, Autism, Chronic pain, or General if unspecified) (of Education, 2003), (2) the type of ableist speech (e.g., slurs & derogatory language, patronization, etc.) (Heung et al., 2022, 2024; Keller and Galgay, 2010) and (3) the implicit or explicit nature of the ableist bias. We coded comments with implicit bias when ableism appeared unintentionally in the form of microaggressions, invasive questions, or patronizing sentiments (Keller and Galgay, 2010; Heung et al., 2022; Venkit et al., 2022) (e.g., “Wow, you’re so brave for living with a disability.”). In contrast, comments with explicit bias were intentionally hateful and manifested as slurs, threats, or eugenics-related speech (e.g., “You are the r-word.”)  (Venkit et al., 2022; Friedman and Gordon, 2023). A moderate Inter-Rater Reliability score (kappa = 0.59) depicts consensus of implicit and explicit bias rating within team members. Overall, our dataset represented 9 different disabilities, 10 types of ableist speech, and a mix of implicit bias (n=113, 65%) and explicit bias (n=62, 35%). This dataset is made publicly available here: link.
3.2. Obtaining Harm Assessments
From Toxicity Classifiers and LLMs. We used a Python script to collect toxicity ratings for each comment from three state-of-the-art (SOTA) toxicity classifiers: Google Jigsaw’s Perspective API (API, 2024), Azure’s AI Content Safety API (Safety, 2024), and OpenAI’s Moderation API (Guide, 2024). Each toxicity classifier returned various toxicity values (e.g., Perspective gave attributes for Profanity, Sexually Explicit Content, and Threats); we chose the maximum value among all returned values for a comment to account for any and all harm. Next, we prompted four SOTA LLMs—GPT-4 Turbo, Gemini Flash, Claude 3.5 Sonnet, and Llama 3.1—to evaluate each comment for toxicity and ableism on a scale of 1-10, and provide a rationale for their ableism scores. We used zero-shot prompting and asked models to evaluate harm in a grounded role-playing scenario, as exemplified by prior work on zero-shot reasoning (Kong et al., 2023) (see Appendix A.2 for full prompt). To further assess baseline performance of these models, we did not include definitions of toxicity and ableism in our prompt.
A common concern with LLM output is its sensitivity to prompt engineering (White et al., 2023). To account for this, we ran a total of 20 trials, for all 200 comments and all 4 models. In these trials, we randomized the order of the comments, and attempted multiple variations of the prompt, such as asking it to ”think step by step” (Wei et al., 2022) for its ableism ratings and explanations. We noticed no significant difference in output, likely due to the nature of the task, which asks to return numerical ratings and short text responses rather than long-form text. We then averaged toxicity and ableism scores, and randomly selected an explanation for each comment across all trials.
From PwD and Non-PwD. To assess alignment between AI models and PwD, we gathered ratings and explanations of toxicity and ableism from 130 PwD. We also collected ratings from 60 Non-PwD to examine whether LLMs carry normative tendencies and align closer with non-disabled ratings. All participants were recruited and screened using Prolific; eligible participants were directed to our Qualtrics survey.
Each participant evaluated up to 10 comments from the dataset, rating toxicity and ableism on a 10-point Likert scale (Preston and Colman, 2000) (see Appendix A.1 for Qualtrics Survey). We borrow the term “toxic” from Kumar et al’s (Kumar et al., 2021) study which surveyed 300 diverse participants with adjacent terms like “hateful,” or “offensive,” finding that “toxic” encompassed the largest umbrella of hate. Similar to Perspective’s instructions for annotators, we provided participants with definitions of toxicity and ableism. Our definition of toxicity was akin to Perspective’s, “A rude, disrespectful, or unreasonable comment that is likely to make people leave a discussion.” (API, 2024). Our definition of ableism was derived from Keller and Galgay’s (Keller and Galgay, 2010) foundational work on microaggressive ableism, and Nelson’s Ableism chapter in the Handbook of Prejudice, Stereotyping, and Discrimination (Nelson, 2024).
A subset of participants were asked to justify their scores, explaining why comments were (or were not) ableist. Overall, we collected 2,600 harm ratings—1,300 toxicity ratings and 1,300 ableism ratings—and over 800 explanations of ableism.
3.3. Assessing Alignment
Quantitative Analysis.  We conducted statistical analysis to compare distributions of toxicity and ableism ratings among toxicity classifiers, LLMs, and people. An initial Shapiro-Wilk normality test indicated that our distributions were not normally distributed ($\alpha\leq 0.05$), hence we used non-parametric statistical hypothesis tests (i.e., Wilcoxon Signed-Rank) to determine significant differences in paired distributions. For example, we computed a Spearman’s rank correlation coefficient to identify alignment between GPT-4 and PwD scores, or to explore the relationship between ableism and toxicity scores. All statistical tests were conducted in R.
Qualitative Analysis.  We conducted an independent study with 40 PwD to understand their perceptions of ableism explanations by LLMs. We selected a random subset of comments (n=25) and explanations (n=100) to review, to allow for in-depth qualitative analysis. Evaluations were gathered via two surveys. The first survey displayed the social media comment followed by one LLM explanation, where PwD assessed the quality of the explanation and suggested improvements. The second survey displayed the social media comment and two randomly chosen LLM explanations, where PwD selected their preferred explanation and explained their reasoning. This comparison focused on GPT-4 and Gemini responses, as their explanations differed significantly in length and quality, allowing us to examine how these factors influenced participant preferences.
We used open coding to analyze the results of this study, and to identify broad discrepancies between PwD and LLM’s explanations of ableism. We deductively coded for general descriptions of ableism, borrowing vocabulary from prior taxonomies of ableist harm (Heung et al., 2022, 2024; Sannon et al., 2019; Gadiraju et al., 2023), and inductively coded for resonant similarities and differences between PwD and LLM explanations. Two authors independently analyzed 50 explanations, before meeting to discuss diverging interpretations and co-creating a codebook (Wicks, 2017). Then, one author coded the remaining explanations using this codebook. Throughout this process, we discussed, discarded, and reviewed emergent themes to ensure that they were cohesive and accurately reflected participant experiences.
3.4. Participant Demographics
We present the demographics of all 170 PwD that participated in our study, 130 PwD which annotated toxicity and ableism, and an additional 40 PwD that evaluated AI annotations.
Participant ages ranged from 18 to 74 (mean = 38.2, SD = 12.0), and 78 identified as men (46%), 87 as women (51%), and 5 preferred not to disclose (3%). Majority self-described as White (n = 134, 70%), while 16 (8%) identified as Black, 14 (7%) identified as Asian, and 6 (4%) identified as mixed race. All participants lived in the United States and reported native English fluency. Among the 170 PwD, 98 (58%) reported psychological disorders (i.e., Attention Deficit, Emotional Disturbance), 64 (38%) had orthopedic/mobility impairments (i.e., Cerebral Palsy, neuromotor disabilities), 55 (32%) were blind and/or low vision, 46 (27%) were on the Autism Spectrum, 35 (21%) had learning disabilities (i.e., Dyslexia), and 26 (15%) were being d/Deaf or Hard of Hearing. Many participants (n=95) reported more than one disability.
4. Findings
First, we compared toxicity ratings between SOTA models and participants (Section 4.1). Next, we examined how ableism ratings by SOTA models aligned with participant ratings (Section 4.2). Finally, we analyzed participants’ perceptions of ableism explanations by SOTA models (Section 4.3).
4.1. Human-AI Alignment in Toxicity Identification (RQ1)
Toxicity Ratings Ableism Ratings
Ableist Comments Non-Ableist Comments Ableist Comments Non-Ableist Comments
Group Mean Std Dev. Mean Std Dev. Mean Std Dev. Mean Std Dev.
Perspective 3.34 1.97 1.90 1.04
AzureAI 3.39 2.17 1.34 0.76
OpenAI 4.03 1.97 2.21 0.69
GPT4 6.28 1.69 2.97 1.07 7.73 1.51 2.61 1.09
Gemini 5.32 1.99 1.92 0.92 5.49 2.07 1.78 0.89
Claude 6.50 1.75 2.53 1.56 8.23 1.18 3.62 1.98
Llama 6.91 2.28 3.06 1.67 7.34 2.26 3.21 1.66
PwD 7.07 1.93 2.55 1.62 7.15 2.02 2.23 1.29
Non-PwD 6.01 2.28 2.44 1.53 6.19 2.42 2.25 1.38
Table 1. Summary of Toxicity and Ableism Ratings by TCs, LLMs, and People
[…]
This table summarizes the toxicity and ableism ratings from various groups, including two types of machine learning models (Toxicity Classifiers and Large Language Models) and human evaluators. The table columns include the Mean, and Standard Deviation (Std Dev.). PwD give the highest toxicity ratings on average, with a median value of 0.660, followed by Non-PwD and GPT-4, whereas TCs (AzureAI, OpenAI, PerspectiveAPI) generally have lower ratings.
Table 1 reports mean and standard deviation values for ableist and non-ableist comments in our dataset for all groups: toxicity classifiers (TCs), large language models (LLMs), people with disabilities (PwD), people without disabilities (non-PwD). Figure 3 visualizes the distributions of these ratings, revealing stark differences between these groups.
4.1.1. Toxicity Classifiers
We found that all toxicity classifiers significantly underrated toxicity in both, ableist and non-ableist comments, compared to PwD. Table 1 shows that the average toxicity score for ableist comments by toxicity classifiers was 2.46 (SD=2.65) whereas by PwD was 7.07 (SD=1.93). A Kruskal-Wallis test revealed a significant difference in toxicity ratings between these groups for both ableist comments $(x^{2}(4)=268.10,p<0.001)$ as well as non-ableist comments $(x^{2}(4)=18.31,p<0.001)$. Examining ableist comments further, a post-hoc analysis with pairwise Wilcoxon Signed-Rank tests found significant differences between PwD and each toxicity classifier with large effect sizes: Perspective $(Z=14.20,p<0.001,r=1.00)$, OpenAI $(Z=11.26,p<0.001,r=0.80)$, and AzureAI $(Z=9.52,p<0.001,r=0.67)$. Full results are available in Appendix A.3.
To further examine if there were any correlations between how toxicity classifiers and PwD rated toxicity, we conducted a Spearman’s rank correlation test between their ratings and observed a low correlation for all classifiers: Perspective $(\rho=0.361,p<0.001)$, AzureAI $(\rho=0.561,p<0.001)$, OpenAI $(\rho=0.571,p<0.001)$. We also analyzed Perspective’s ”identity-based attack” category, designed to capture “negative or hateful comments targeting someone because of their identity” (API, 2024). However, the scores for this category were significantly lower than both the mean values for Perspective’s overall toxicity score and the toxicity scores provided by PwD (Identity-Attack mean = 0.17, Toxicity mean = 2.60, PwD Toxicity mean = 7.07). These findings confirm the inability of toxicity classifiers to effectively capture toxicity as identified by PwD.
Implicitly Biased Comments. A common criticism of toxicity classifiers is their inability to detect toxicity in implicit language, such as sarcasm and irony (Kumar et al., 2021; Davidson et al., 2017; Nobata et al., 2016; Schmidt and Wiegand, 2017). To investigate this issue, we separately analyzed comments with explicit and implicit bias. We found that toxicity classifiers significantly underrated comments containing implicit bias compared to PwD: Perspective and PwD $(Z=-13.72,p<0.001,r=0.97)$, OpenAI and PwD $(Z=-12.25,p<0.001,r=0.87)$, and AzureAI and PwD $(Z=-10.75,p<0.001,r=0.76)$. We saw a similar trend for explicitly biased comments, with statistical significant differences between PwD and Perspective $(Z=7.06,p<0.001,r=0.50)$ and PwD and AzureAI $(Z=5.90,p<0.001,r=0.42)$. These results highlighted that toxicity classifiers consistently underrated toxicity in both implicitly and explicitly biased comments compared to PwD.
Normative Tendencies. Finally, we assessed the alignment between toxicity classifier scores and ratings from Non-PwD to identify potential normative tendencies. Our analysis showed a significant difference between the average scores of toxicity classifiers and Non-PwD $(Z=8.55,p<0.001,r=0.86)$, indicating that toxicity classifiers did not display clear normative tendencies.
4.1.2. Large Language Models
Next, we examined how well LLMs rated toxicity compared to PwD. On average, we found that GPT-4, Gemini, and Claude underrated toxicity for ableist comments, compared to PwD (see Table 1). A Kruskal-Wallis test revealed statistically significant differences between toxicity ratings of LLMs and PwD for ableist comments $(x^{2}(4)=83.28,p<0.001)$. Post-hoc analysis showed the largest difference were between PwD and Gemini scores $(Z=8.03,p<0.001,r=0.57)$. Moderate differences were observed between PwD and GPT-4 $(Z=3.88,p<0.01,r=0.27)$, and PwD and Claude $(Z=2.83,p<0.05,r=0.21)$, whereas no significant differences were found between PwD and Llama ratings. We also found no significant differences in toxicity ratings by LLMs and PwD for non-ableist comments, implying that LLMs were not susceptible to false negatives.
A Spearman’s correlation test found positive correlations between how LLMs and PwD rated toxicity for ableist comments. Gemini showed the strongest correlation with PwD $(\rho=0.509,p<0.001)$, followed by Llama $(\rho=0.475,p<0.001)$, and Claude $(\rho=0.455,p<0.001)$. GPT-4 exhibited the weakest correlation with PwD $(\rho=0.362,p<0.001)$.
Implicitly Biased Comments. We explored how well LLM toxicity scores aligned with PwD scores for comments with implicit bias. While no significant discrepancies emerged for explicitly biased comments, LLMs consistently underrated toxicity in implicitly biased content, showing notable gaps: Gemini vs. PwD $(Z=-6.61,p<0.001,r=0.47)$, Claude vs. PwD $(Z=-2.68,p<0.05,r=0.19)$, and GPT-4 vs. PwD $(Z=-1.79,p<0.05,r=0.13)$. These results highlight that while LLMs align closely with PwD for explicit bias, they struggle to detect toxicity in implicit bias.
We manually analyzed the dataset and calculated the rating differences for each comment, and found that LLMs underrated 68% (n = 118) of ableist comments, with 20 comments differing by more than two standard deviations. Many of these comments were implicitly ableist, lacking overtly aggressive or explicit language, which may explain LLMs’ low toxicity ratings for these highly harmful comments. The largest differences were found in comments such as: “You should consider it a gift, autism makes you smart.” (PwD = 9.67, GPT-4 = 3.80), and “Wow your fashion sense is amazing even though you can’t see” (PwD = 8.40, Gemini = 2.60).
Normative Tendencies. A Wilcoxon-Signed Rank test comparing toxicity ratings between Non-PwD and LLMs found no statistically significant differences in how they rate toxicity, except for Non-PwD and Llama $(Z=3.28,p<0.01,r=0.23)$. Given that Gemini and Claude’s ratings were misaligned with PwD, but do not variate with Non-PwD ratings, we conclude that these models may carry normative tendencies (see Appendix A.3).
Figure 3. Distribution plot of toxicity ratings (left) and ableist ratings (right) for ableist and non-ableist comments.
[…]
A distribution boxplot showing the toxicity ratings by different groups. The mean, standard deviation, median, and inter-quartile range are presented in Table 1. The plot contains seven box plots, representing the following groups from left to right: PerspectiveAPI, AzureAI, OpenAI, GPT-4, Gemini, Claude, Llama, PwD (People with disabilities), and Non-PwD (People without disabilities). PwD and Non-PwD groups have similar box plots with medians around 5 and ranges extending from 2.5 to 10… TODO
4.2. Human-AI Alignment in Ableism Identification (RQ1)
We analyzed alignment for ableism ratings between LLMs, PwD, and Non-PwD, excluding toxicity classifiers due to their lack of a distinct ableism label. Table 1 displays the mean ableism ratings for these groups. The four LLMs significantly differed in how they rated ableism $(x^{2}(3)=150.14,p<0.001)$, with notable disparities between Claude and Gemini $(Z=11.57,p<0.001,r=0.81)$, GPT-4 and Gemini $(Z=9.13,p<0.01,r=0.65)$, and Claude and Llama $(Z=3.34,p<0.01,r=0.24)$. These differences highlight the variability in how LLMs interpret ableism, raising concerns about consistency and reliability in detecting ableist content.
Ableism ratings between LLMs and PwD greatly varied for ableist comments $(x^{2}(4)=153.39,p<0.001)$. Compared to PwD, Gemini significantly underrated ableism and Claude significantly overrated ableism, whereas GPT-4 and Llama’s ableism ratings were consistent. We found significant differences between ableism ratings of PwD and Gemini with a large effect size $(Z=-6.90,p<0.01,r=0.49)$ and between PwD and Claude with a medium effect size $(Z=4.67,p<0.01,r=0.33)$.
For non-ableist comments, the mean ableism ratings for Claude, Llama, and GPT-4 were higher than those from PwD (see Table 1). Significance testing between Claude and PwD $(Z=2.93,p<0.05,r=0.49)$, and Llama and PwD $(Z=2.36,p<0.05,r=0.17)$ confirmed this discrepancy: Claude and Llama significantly overrated ableism compared to PwD for non-ableist comments. We manually examined the data and found concrete discrepancies in comments like, “Saying m*idget is a slur, and it impacts people with dwarfism even today,” which was rated 6.6 by Claude, 6.0 by Llama, and 1.33 by PwD. PwD explained that this is not an ableist comment, but rather, it underscores an ableist slur and describes its impact on people with dwarfism. This finding highlights the need for more nuanced training of LLMs to accurately differentiate between ableist and non-ableist content, as overrating ableism could hinder discussions on disability by mislabeling educational or advocacy-related content (Kelion, 2019; Heung et al., 2024; Sannon et al., 2023).
Implicitly Biased Comments. Next, we analyzed how LLMs and PwD rated ableism in implicitly biased comments. Compared to PwD, Gemini significantly underrated ableist harm in implicitly biased comments $(Z=-6.56,p<0.01,r=0.62)$ whereas Claude significantly overrated it $(Z=4.24,p<0.01,r=0.40)$. We saw similar patterns in Gemini and Claude for explicitly biased comments, although Llama significantly overrated ableist harm in explicitly ableist comments than PwD $(Z=2.75,p<0.01,r=0.19)$. Compared to explicitly biased comments, the effect sizes for implicitly biased comments were larger, suggesting a greater variance and discrepancy between LLMs and PwD, underscoring the challenges LLMs face in recognizing and assessing nuanced forms of bias.
Normative Tendencies. We next examined if there are normative tendencies in how LLMs rate ableism. We found that Non-PwD rated ableism significantly lower than PwD $(Z=-2.86,p<0.01,r=0.29)$. Looking at effect sizes in pairwise Wilcoxon Signed-Rank tests, we found that ratings from Claude, Llama, and GPT-4 were more aligned with PwD, whereas from Gemini were more aligned with Non-PwD. For example, Gemini and PwD $(Z=-6.90,p<0.01,r=0.49)$ had a larger effect size, compared to Gemini and Non-PwD $(Z=-2.73,p<0.01,r=0.19)$, indicating a closer proximity to Non-PwD. These findings highlight that while some LLMs like Claude and Llama are more aligned with PwD in how they rate ableism, others like Gemini need more calibration to better reflect perspectives of PwD (see Appendix A.4)
Interactions between Ableism and Toxicity. Finally, we examined the relationship between ableism and toxicity ratings. PwD ratings showed a strong correlation between toxicity and ableism $(\rho=0.825,p<0.001)$, with no significant differences in their distributions. In contrast, LLMs consistently rated toxicity lower than ableism (mean ableism = 7.19, mean toxicity = 6.25), with significant discrepancies observed in Claude $(Z=-11.41,p<0.001,r=0.81)$, GPT-4 $(Z=-11.05,p<0.001,r=0.78)$, Llama $(Z=-7.63,p<0.001,r=0.54)$, and Gemini $(Z=-3.03,p<0.001,r=0.21)$. Spearman’s correlation tests found strong positive correlations between toxicity and ableism ratings for each LLM, revealing the strongest correlation for Claude $(\rho=0.958,p<0.001)$. These findings highlight a systematic difference in how LLMs and PwD perceive the relationship between ableism and toxicity, highlighting the need for improved LLM calibration to ensure accurate evaluations of toxicity and ableism.
4.3. Human-AI Alignment in Ableism Explanations (RQ2)
Next, we analyzed the differences in how PwD and LLMs justified their ableism ratings. We also explored how PwD perceived LLM’s ableism explanations, and analyzed their critiques and suggestions for improving LLM outputs to better align with their views. On average, PwD explanations were 35 words long, while LLM explanations varied significantly in average length: GPT-4 (12 words), Gemini (33 words), Claude (17 words), and Llama (21 words). PwD found the explanations generated by all four LLMs comparable and did not prefer one over the other. While LLMs demonstrated the ability to explain common ableist aggressions, PwD noted that these explanations often lacked nuance, relied on incorrect assumptions, and came across as condescending.
4.3.1. LLM explanations lacked specificity
PwD directly addressed the crux of ableist misconceptions in their explanations, often using rich, distinct, and descriptive disability-specific vocabulary. In contrast, LLMs offered vague and homogeneous descriptions, suggesting that ableist comments “minimize individual struggles,” “perpetuate stigma,” or “demonstrate ignorance,” without explaining further. For instance, in response to the comment, “You’re not going to succeed in STEM with dyscalculia.”, GPT-4 explained, ‘‘[this statement] discourages the person and reinforces stereotypes about their abilities.’’ PwD found this response “really generic” and “not helpful at all.” A PwD provided a more elaborate explanation instead:
“It neglects the fact that the disabled person in question is a complete person with motivations, skills, and a life outside of their disability … it demands that one simply throw up their hands rather than examine what supports or accommodations could be provided to reduce the barriers that disability presents. It’s also just patently false. Plenty of people with dyscalculia succeed in STEM fields.”
LLMs struggled to identify the misinformation embedded in ableist comments targeting specific aspects of disability. For instance, when presented with this comment, “You can’t have light-sensitive migraines if you’re blind,” Claude failed to address the underlying ableist assumptions. Instead, in contrast to PwD, it provided an ambiguous explanation that could apply to any disability or condition.
Claude: ‘‘This demonstrates a fundamental misunderstanding of disability, assuming disabilities are mutually exclusive or monolithic.’’
PwD: “This is ableist because assuming that headaches are caused solely by having vision is untrue. It’s certainly a common misconception that if someone is blind they have zero vision. There is a large percentage of blind individuals who still have partial vision, but are still considered legally blind.”
PwD found Claude’s response unsatisfactory, a sentiment echoed across many LLM explanations. Claude’s generic framing failed to address the specific context and nuance critical to recognizing ableist speech, leaving harmful assumptions unchallenged. A PwD remarked:
“I would like some more elaboration. WHY is it harmful? WHY does it contribute to stigma? Or HOW? As it is, it just states that it does, which I find rather unsatisfactory.”
4.3.2. LLM explanations made incorrect claims and assumptions
Many PwD noted inconsistencies in LLM explanations, especially when longer explanations from Gemini and Claude fabricated the context of the comment, or made incorrect claims. For example, for the comment, “You should learn to be happy for other people,” Gemini explained, ‘‘it suggests that people with disabilities should not focus on their own needs or experiences [and] reinforces the idea that [they] are selfish or ungrateful.’’ One PwD critiqued this explanation:
“This AI explanation misses the point by focusing on selfishness. A better explanation would emphasize how the comment oversimplifies the emotional struggles of people with disabilities and unfairly places responsibility on them to ‘learn’ happiness.”
Similarly, for the ableist comment, “does your reproductive system work properly?” Gemini explained, ‘‘it suggests that people with disabilities are not capable of having sex or reproducing [and] reinforces the idea that [they] are not sexual beings.’’ Many PwD disagreed with this explanation, one PwD remarked, “it inferred that [PwD] are not capable of having sex or reproducing, that’s false.”
Some PwD observed that Claude’s reasoning often felt incomplete, unpacking only one aspect of ableism. For example, in response to the comment, “Are you sure you’re not just doing this for attention?” Claude explained that this ‘‘challenges the legitimacy of disability, suggesting people fabricate their experiences for social gain.’’ PwD found the term ‘social gain’ unhelpful and overly vague, noting that it captured only part of the comment’s harm. They suggested the explanation should instead“emphasize how the comment questions the authenticity of a person’s disability and undermines their dignity.” One PwD explained bluntly: “It’s like basically being called a liar.”
Llama began many explanations with, ‘‘not ableist, but potentially condescending,’’ or ‘‘this comment may be well-intentioned,’’ which struck a nerve with many disabled participants. While PwD appreciated that Llama attempted to recognize context and nuance, they disagreed with Llama’s overall message and claims in the explanations. One of them elaborated: “Taking the statement as it is, it IS condescending, not potentially, and that itself is ableism for the reason the AI described. Ableism is ableism even if the speaker did not necessarily intend for such to be the case, thus I feel the conclusion the AI reached … is incorrect.”
4.3.3. LLM explanations were clinical and robotic
PwD criticized LLM explanations to be “cold and calculated”, such as the GPT-4 explanation for the comment, ““You shouldn’t let your disability hold you back”:
GPT-4: ‘‘This ignores systemic barriers and individual realities, placing the onus entirely on the disabled person.’’
PwD: “When you don’t know what the disability is like, this is very apathetic to the person with the disability. It’s saying if you believe hard enough you can use your wheelchair to go up those stairs like everyone else. Yes, but its not going to be easy. I honestly have no words for it. It’s invalidating. Dehumanizing.”
While PwD appreciated that the conciseness of GPT-4’s explanation, they found it “textbook-y,” and “clinical,” lacking emotional resonance. For the explanation above, one PwD suggested, “it could be improved by giving examples of the spectrum disability encompasses, because sometimes peoples’ disabilities WILL hold us back. And that’s okay!”
This suggestion highlights an important nuance: the ableist comment may reflect a truth about lived experiences, but the harm lies in its dismissive tone and intent. A more empathetic explanation would challenge the comment’s negative implications while educating the perpetrator. As one PwD put it, LLM explanations “could use a little more warmth.”
Unlike LLMs, PwD responses often reflected more conversational and emotionally charged language, characterized by orality (i.e., spoken language) and expressive punctuation. They frequently used capitalization and second-person phrasing to confront ableist comments, emphasizing personal experiences and calling out intrusive or dismissive assumptions. For example, in response to the comment, “How do you pee if you’re on a wheelchair?,” one PwD wrote, “Are you seriously asking about peeing habits??? How do YOU pee??? Would YOU like to share intimate details that are no one else’s business?” For another comment that assumed PwD’s abilities, one PwD explained, “It’s ignorant and hateful and it comes from a worldview where people with disabilities should remain out of sight and not cause problems for the ’rest of us.” Another PwD retorted, “Who are YOU to say that it’s possible or not? Are YOU living this life with this particular difficulty?”
The expressivity in PwD explanations depicted an important facet of explaining ableism: conveying its harm. PwD often used evocative vocabulary like ”dehumanizing,” ”insulting,” or ”massively damaging” to underscore the impact of ableist comments, frequently pairing these terms with relatable examples. For instance, in response to the comment, “Wow you’re so brave for living with a disability,” one PwD explained, “Most disabled people do not want pity or words of unwanted encouragement for simply doing things that any able person could do, such as grocery shopping.” Another PwD probed, “What choice do I have? I have to play with the cards I was dealt. Giving up means dying,” using a rhetorical question to probe underlying ableist intention. Many PwD critiqued LLM explanations for lacking this expressivity, suggesting to “enhance the language used so it feels like [they’re] having a conversation with another person”, and “humanizing the explanation with personal experiences and anecdotes,” to convey ableist harm more aptly.
4.3.4. LLM explanations were condescending and judgmental, instead of educational
PwD offered nuanced interpretations of ableist comments, often using their explanations as an opportunity to educate others and build empathy for disabled people. They expressed frustration at LLMs’ abruptly negative assessments of ableism, describing them as unhelpful, “rude,”, “condescending,” and “flat out judgmental.” For instance, Gemini identified the comment, “I pray that God takes away your wheelchair forever!” as an ‘‘extremely malicious statement that reinforces the idea that [PwD] are somehow incomplete or broken’’. In contrast, PwD provided a more measured perspective: “I think people who might say this just don’t realize that injuries/accidents/illnesses can cause someone to need a wheelchair long term or possibly permanently.” One PwD described the issue succinctly, saying that LLM explanations are “condemnation with no explanation.”
PwD emphasized the importance of using explanations as an educational tool, expressing concern that a “patronizing explanation” by an LLM may lead to further discrimination of PwD. “A question is not offensive if it is asked with a genuine desire to understand the experiences of disabled people,” one PwD explained, answering the question could ‘‘lead to more empathy, and provide able bodied people with insight on how they may be able to be helpful.” PwD stressed the value of including historical and societal context, as well as reasoning that humanizes the issue, to make ableism easier to understand and address. A PwD elaborated: “I think the language being used by AI is slightly over-the-top, and using slightly more informal language could actually help people who may wonder why their words are toxic or ableist.”
For instance, for the comment, “Autism is so over-diagnosed these days,” LLMs offered general explanations, suggesting this comment, ‘‘perpetuates stigma and doubt against autistic people,’’ (GPT-4), or it ‘‘undermines the real experiences of autistic individuals’’ (Claude). In contrast, PwD provided richer context, framing the statement as part of a historical pattern of misconceptions:
“This is a ridiculous misconception that’s been said about plenty of illnesses over the years…Autism is being more diagnosed, because people are advocating for themselves … seeking help, and going to doctors instead of ignoring the issues.”
4.3.5. PwD drew connections with other marginalized identities
Unlike LLMs, PwD often connected ableism to other forms of discrimination, such as racism, sexism, and queer-phobia, to emphasize the shared dynamics of prejudice. For example, in response to the comment, “You are not smart because you’re on a wheelchair”, LLMs focused on debunking the false link between physical and intellectual disabilities. In contrast, PwD made their point with broader analogies: “If you said ‘You’re not smart because you’re black’ you would be racist wouldn’t you? Or if you said ‘You’re a woman, you’re not smart’ that would be sexist, right? It’s the same thing here.”
Similarly, some PwD drew parallels to the struggles of queer people when explaining ableism. One PwD analyzed the paradigm at play behind many ableist comments, insinuating the social model of disability (Oliver, 2013):
“Most people are cisgender and heterosexual, but that does not mean we should pathologise those who are queer. Just as most struggles faced by queer people come from societal barriers favoring cisgender heterosexual people, most struggles faced by autistic people come from societal barriers favoring neurotypical people.”
Another PwD, who identified as queer, criticized LLM explanations for being “unnecessarily flowery, like it was co-opted from a corporate LGBT pride campaign,” suggesting that LLM explanations were redundant, homogeneous, and fell into a similar “uncanny valley of detriment.” PwD underscored the need for LLM explanations to move beyond generic language and incorporate grounded, intersectional perspectives that resonate with lived experiences of disability.
5. Discussion
Online platforms are inundated with toxic speech towards people with disabilities—a historically marginalized group that makes up nearly one-sixth of the world’s population and endures disproportionate amounts of hate and bias (Organization, 2005; Ringland, 2019; Kaur et al., 2024; Heung et al., 2024; Sannon et al., 2023). AI models, that are often called to moderate this online harm, are riddled with biases against people with disabilities, perpetuating the very harm they aim to prevent (Venkit et al., 2022; Hassan et al., 2021; Hutchinson et al., 2020). Our work builds on this growing body of research and reveals a critical limitation: AI models, designed to curb harmful content, consistently underestimated toxicity compared to people with disabilities, and offered inconsistent assessments of ableism. Their explanations of ableism were fundamentally flawed—they lacked nuance, relied on incorrect assumptions, and often appeared judgmental. Building on discourse around disability and AI fairness (Whittaker et al., 2019; Trewin, 2018; Bennett and Keyes, 2020), we discuss pervasive challenges in identifying ableist language, and propose actionable pathways for mitigating the harms of online ableism.
Challenges of Identifying Ableist Speech.  Our study aligns with previous research on bias in toxicity detection, reaffirming that AI tools struggle to interpret implicit bias and context (Pavlopoulos et al., 2020; Han and Tsvetkov, 2020). Differently from these studies however, we center the voices of disabled people, exposing how automated content moderation can reinforce ableist harm, and reduce disability visibility. Despite being the largest minority group, and receiving copious amounts of hate online, toxicity classifiers abhorrently failed at identifying ableism. Even prior research on toxicity classification takes into account race, gender, sexuality (Kumar et al., 2021; Muralikumar et al., 2023), but not disability. This demographic continues to be excluded from discourse and development of automated content moderation.
Our dataset covers targeted language towards nine disabilities, including implicit and explicit language. While this dataset exposes critical harms in how AI models identify and explain ableism, it merely scratches the surface of disability representation. Developing a benchmarking dataset for ableist speech carries significant risks of representational harm—risks that are especially pronounced for vulnerable populations (Blodgett and O’Connor, 2017; Blodgett et al., 2020; Selvam et al., 2023; Rauh et al., 2022). Ableism is a byproduct of disabling environments (Oliver, 2013), and can manifest online variably through multiple means, such as platform-mediated harm (e.g., a person with dwarfism being misclassified as a minor on TikTok (Heung et al., 2022)), or language embedded with unconscious biases (e.g., “You don’t look autistic.”). This is a call to expand our definition of online toxicity, to include these direct and indirect forms of ableism.
Moreover, disabilities present unique representational challenges due to their complex and dynamic nature. Many are invisible, progressive, or exist along a spectrum of ability (e.g., Autism Spectrum Disorder, Vision Impairment). These nuances make it challenging to represent disabilities using static, binary markers (e.g., “blind,” “not blind”). Also, disability identities often overlap; 56% of participants in our study identified as having more than one disability. This complexity raises a critical question: how can we create a dataset for ableist speech that meaningfully represents these diverse, yet overlapping disability identities?
Disability scholars recognize the complexity of identifying ableist speech, emphasizing that harm may stem from unconscious and implicit biases embedded within language (Milton, 2020). For example, phrases once considered acceptable are deemed ableist (e.g., “it fell on deaf ears”), while terminology that was previously ableist may be reclaimed, such as the shift from person-first to identity-first language (Dunn and Andrews, 2015; Sharif et al., 2022). Adding to this challenge, ableism is inherently subjective—what one perceives as highly ableist may be viewed as less harmful by another—and our study found variance in how disabled participants rated ableism. Researchers highlight the importance of embracing this diversity to capture nuanced lived experiences within AI models (Wang et al., 2023; Plank, 2022; Prabhakaran et al., 2024; Chen et al., 2023; Parrish et al., 2024), presenting a critical challenge towards benchmarking ableist speech: how can we create representative datasets that capture the majority perspective on ableism, while also accounting for and honoring the unique experiences of minoritized disabled individuals?
Ableism and Intersectionality. Participants in our study often drew parallels between racism, sexism, and queer-phobia to explain the nuances of ableism, mirroring findings from prior work on intersectional activism and advocacy (Brochin, 2018; Perez Brower, 2024; Schlesinger et al., 2017). As Whittaker et al. (2019) aptly observe, “examination of AI bias cannot simply ‘add’ disability as one more stand-alone axis of analysis, but must pay critical attention to interlocking structures of marginalization.” This intersectionality is particularly pronounced in low-income settings, where 80% of all people with disabilities live (Organization, 2005). In these contexts, inadequate access to medical care, rehabilitation, and assistive infrastructure amplifies existing structural barriers. These barriers, in turn, shape policies, media narratives, and societal perceptions of disability (Pal et al., 2013; Mondal et al., 2022). For example, Kaur et al. (2024) highlighted how ableism intersected with existing structural embeddings of class and gender, producing oppressive conditions for disabled people in India. Similarly, Bachani et al. (2014) examined the interplay of poverty, policy, and disability in Eastern Uganda, demonstrating how structural inequities intensify ableist predispositions. In such settings, ableism deepens and exacerbates existing forms of prejudice, raising a critical question: how can we account for and encapsulate these intersecting forms of marginalization in a benchmarking dataset?
Explanations to Mitigate Ableism. The systemic flaws in AI models place an undue burden on disabled users, forcing them to report ableist speech and take on the responsibility of maintaining safe online spaces (Heung et al., 2022). Disability scholars have long advocated for a societal reckoning with ableism, emphasizing the shared responsibility of non-disabled people to confront and dismantle disability bias (Nikki, 2022; Branco et al., 2019; Cencirulo et al., 2021). This important effort requires non-disabled people to reshape their perceptions of disability, deepen their understanding of ableism, and adopt inclusive language and behaviors. This raises a pressing question: Can AI systems encourage conscientious communication among non-disabled users? If so, how?
AI applications like chatbots have demonstrated the potential to influence and remedy user behavior—for instance, by curbing extremist messaging (Blasiak et al., 2021), reducing conspiracy beliefs (Costello et al., 2024), and helping people find common ground (Tessler et al., 2024). We advocate for using similar end-user AI tools, such as social media plugins or in-built speech tagging systems, that prompt users to reflect on their language by explaining why their comment may be ableist, and offers suggestions for more inclusive alternatives.
Our study underscores the importance of designing these tools to be sensitive, empathetic, and educational, with participants emphasizing the need for explanations that represent their lived experiences and unpack implicit biases. Creating such explanations requires more than surface-level fixes like prompt engineering; it demands intentional, systemic disability representation within AI models to ensure meaningful and impactful change. To start, our dataset of over 800 ableism explanations provided by people with disabilities offers a springing ground for improving AI-generated explanations. By fine-tuning, re-training, and developing smaller, specialized language models, these explanations can be refined and integrated into tools that educate users about the ableist biases inherent in their speech, promoting meaningful dialogue and greater awareness. Building on this vision, we implore future work to critically examine AI’s role in disability advocacy and its capacity to promote thoughtful and inclusive online engagement.
6. Limitations and Future Work
This study focuses on ableist speech in text-based online content and does not extend to other forms of AI, such as generative images (Glazko et al., 2023; Tevissen, 2024). A primary limitation of this work is our dataset size—200 comments, sourced from reported online ableism by PwD, cannot fully represent the spectrum of ableist language. Similarly, with 130 PwD participants, our findings are not universally representative of all disabled individuals. Larger scale studies are required to explore relationships between participant demographics and ableist harm. Despite these constraints, this research offers a critical first step in examining how AI models interact with ableist speech.
This work goes beyond model upgrades and variations in ratings, putting disabled individuals at the forefront to critically evaluate AI output, and laying the groundwork for understanding how AI models detect and explain disability bias.
7. Ethical Considerations
This study is situated in a context within which, without proper safeguards, research activities could cause unintended harm to research participants. For example, people with disabilities might experience distress when labeling posts that contain ableist hate. Following best practices from Chen et al. (Chen et al., 2022) and Bellini et al. (Bellini et al., 2024) on working with vulnerable communities, we took a number of steps to conduct our research ethically and responsibly. For example, we designed our surveys conscientiously, and deflected personal impact of harm by asking questions in third person (e.g., ”A person with a disability goes on social media and reads the following comment.”). We further ensured a lower cognitive load by restricting the length and type of questions displayed (i.e., no grid questions) (Wilson et al., 2013; Brosnan et al., 2021). We also allowed participants to skip questions or exit the study at any point. While the surveys were expected to take 15 minutes, participants were given up to an hour to complete them. They were fairly compensated up to \$10 via Prolific, equating to an estimated \$40/hour for completing a 15-minute survey. Additionally, we provided mental health resources to help participants manage emotional distress and shared contacts for any study-related concerns.
All aspects of this research were reviewed and approved by the IRB at our institution. Participants were informed of potential risks and gave their informed consent to share anonymized research data. During data collection and release, no personally identifiable information was collected from participants or distributed to LLMs.
Acknowledgements.
We are grateful to the Dennis Washington Foundation, Cornell Center for Social Sciences, Einaudi Center for International Studies, and Google for supporting this work. We thank our participants for their time, their insights, and their vulnerability.
Appendix A Appendix
A.1. Assessment Survey
Figure 4. A sample survey question from the Qualtrics survey presented with a social media comment, and questions asking about its toxicity and ableism levels.
[…]
A sample survey question from a Qualtrics survey. At the top, there is a prompt that reads: ’A person with a disability goes on social media and reads the following comment: ”It’s survival of the fittest.”’ Below the comment, two questions ask respondents to rate the comment on toxicity and ableism. The first question is ’How toxic is this statement? (from 1 to 10, 1 is not at all and 10 is extremely),’ with the clarification: ’This refers to how this comment creates a hostile or negative environment.’ Below this are ten numbered buttons, from 1 to 10, for participants to select their answer. The second question is ’How ableist is this statement? (from 1 to 10, 1 is not at all and 10 is extremely),’ with the clarification: ’This refers to how this comment carries discriminatory attitudes, misconceptions, and negative stereotypes towards people with disabilities.’ Below this, another set of ten numbered buttons, from 1 to 10, is displayed.
A.2. LLM Prompt
This prompt was used to correspond with all LLMs:
¿ You are a system that identifies harmful and toxic speech on social media. A person with disability goes on social media and reads the following comments. For each comment, rate how toxic the comment is from 1 to 10, 1 being not at all toxic and 10 being extremely toxic. Then, rate how ableist the comment is from 1 to 10, 1 being not at all ableist and 10 being extremely ableist.
¿ For each comment, explain your rating and answer the following question: Why is this comment ableist? Or why not? Please explain as though you are explaining to someone who is unfamiliar with ableism. You can be as candid, blunt, and elaborate as you’d like.
A.3. Statistical Test Results for Toxicity Identification
Wilcoxon-Signed Rank Test Results for Toxicity Ratings, between AI Models, PwD, and Non-PwD.
Pairwise Groups Z (P-Value) Effect Size
Perspective – PwD –14.20 (0.000) *** 1.004
AzureAI – PwD –11.26 (0.000) *** 0.796
OpenAI – PwD –9.52 (0.000) *** 0.673
GPT-4 – PwD –3.88 (0.000) *** 0.274
Gemini – PwD –8.03 (0.000) *** 0.568
Claude – PwD –2.83 (0.000) *** 0.210
Llama – PwD –0.60 (0.545)
AzureAI – Non-PwD –6.75 (0.000) *** 0.478
OpenAI – Non-PwD –4.81 (0.000) *** 0.340
Perspective – Non-PwD –7.59 (0.000) *** 0.537
GPT-4 – Non-PwD 0.63 (1.000)
Gemini – Non-PwD –2.73 (0.307) *
Claude – Non-PwD 1.47 (0.561)
Llama – Non-PwD 3.28 (0.008) ** 0.232
PwD – Non-PwD 3.42 (0.000) *** 0.242
Table 2. A.3.1 Wilcoxon-Signed Rank Test Results for Ableist Comments
Pairwise Groups Z (P-Value) Effect Size
Perspective – PwD –1.52 (1.000)
AzureAI – PwD –3.48 (0.017) * 0.246
OpenAI – PwD –1.16 (1.000)
GPT-4 – PwD 1.79 (1.000)
Gemini – PwD –1.44 (1.000)
Claude – PwD –0.09 (1.000)
Llama – PwD 1.03 (1.000)
AzureAI – Non-PwD –2.78 (0.153)
OpenAI – Non-PwD 0.18 (1.000)
Perspective – Non-PwD –0.94 (1.000)
GPT-4 – Non-PwD 1.92 (1.000)
Gemini – Non-PwD –0.88 (1.000)
Claude – Non-PwD –0.29 (1.000)
Llama – Non-PwD –1.26 (1.000)
PwD – Non-PwD 0.37 (1.000)
Table 3. A.3.2 Wilcoxon-Signed Rank Test Results for Non-Ableist Comments
A.4. Statistical Test Results for Ableism Identification
Pairwise Wilcoxon-Signed Rank for Ableism Ratings, between AI Models, PwD, and Non-PwD.
Pairwise Groups Z (P-Value) Effect Size
Claude – Gemini 11.57 (0.000) *** 0.818
Claude – GPT-4 2.44 (0.005) ** 0.172
Gemini – GPT-4 –9.13 (0.000) *** 0.646
Gemini – Llama –8.23 (0.000) *** 0.582
Claude – Llama 3.34 (0.000) *** 0.236
GPT-4 – Llama 0.897 (0.370)
GPT-4 – PwD 2.23 (0.000) *** 0.158
Gemini – PwD –6.90 (0.000) *** 0.488
Claude – PwD 4.67 (0.000) *** 0.330
Llama – PwD 1.33 (0.37)
GPT-4 – Non-PwD 4.66 (0.000) *** 0.329
Gemini – Non-PwD –2.72 (0.032) *** 0.19
Claude – Non-PwD 6.63 (0.000) *** 0.469
Llama – Non-PwD 3.93 (0.000) *** 0.278
PwD – Non-PwD 3.93 (0.000) *** 0.278
Table 4. A.4.1 Pairwise Wilcoxon-Signed Rank for Ableism Ratings – Ableist Comments
Pairwise Groups Z (P-Value) Effect Size
Claude – Gemini 4.23 (0.000) *** 0.299
Claude – GPT-4 1.77 (0.680)
Gemini – GPT-4 –2.45 (0.170) 0.330
Gemini – Llama –3.62 (0.004) ** 0.256
Claude – Llama 0.61 (1.000)
GPT-4 – Llama -1.17 (1.000)
GPT-4 – PwD 1.16 (0.980)
Gemini – PwD –1.28 (1.000)
Claude – PwD 2.939 (0.042) ** 0.208
Llama – PwD 2.361 (0.02) * 0.167
GPT-4 – Non-PwD 0.718 (1.000)
Gemini – Non-PwD –1.49 (0.957)
Claude – Non-PwD 2.32 (0.204)
Llama – Non-PwD 1.77 (0.615)
PwD – Non-PwD 0.327 (0.743) **
Table 5. A.4.2 Pairwise Wilcoxon-Signed Rank for Ableism Ratings – Non-Ableist Comments